OpenStudy (anonymous):

(mx - md + b - c)^2 + x - v - r^2 = 0 solve for x

OpenStudy (anonymous):

(mx - md + b - c)^2 + x - v - r^2 = 0 m^2x^2 -2m^2xd +2mxb - 2mxc + m^2d^2 - 2mdb + 2mdc +b^2 -2bc +c^2 + x - v - r^2 = 0 x^2(m^2) + x(-2m^2d+2mb-2mc+1) + m^2d^2 - 2mdb + 2mdc +b^2 -2bc +c^2 - v - r^2 = 0 a = m^2 b = -2m^2d+2mb-2mc+1 c = m^2d^2 - 2mdb + 2mdc +b^2 -2bc +c^2 - v - r^2 x = (-b±sqrt(b^2-4ac))/2a right?